OpenStudy (earth04):

sqaure root of (2x+6) minus square root of (x+4)=-1

OpenStudy (anonymous):

\[\sqrt{2x+6}-\sqrt{x+4}=-1\] To ease up on the calculations, re-order like so: \[\sqrt{2x+6}+1=\sqrt{x+4}\] Raise both sides to the power of 2: \[\left( \sqrt{2x+6}+1 \right)^2=x+4\] Try to continue from here.